OpenStudy (anonymous):

Write the equation of the line that passes through (–4, 3) and is perpendicular to the line y=7/5x-3/5. A. 5x + 7y = 1 B. 7x + 5y = –13 C. 7x + 5y = –1 D. 3x + 5y = 7

OpenStudy (anonymous):

you know the form of equation y = mx + c where m is the slope of line and c is some constant?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

here the slope of given line is y = mx+c , we need m value so, m =7/5 and slopes of perpendicular lines have a property if m1 and m2 are slopes of perpendicular lines then product of slopes = -1 so m*7/5 = -1 => m = -5/7 { this is what the slope of our required line

OpenStudy (anonymous):

now you have a point (–4, 3) and slope -5/7, use point-slope formula :)

OpenStudy (anonymous):

ohh ok I get it

OpenStudy (anonymous):

:)